### v1

# DRESS: A Continuous Framework for Structural Graph Refinement

Abstract
The Weisfeiler-Lehman (WL) hierarchy is a cornerstone framework for graph isomorphism testing and structural analysis. However, scaling beyond 1-WL to 3-WL and higher requires tensor-based operations that scale as $\mathcal{O}(n^{3})$ or $\mathcal{O}(n^{4})$, making them computationally prohibitive for large graphs. In this paper, we start from the Original-DRESS equation (Castrillo, León, and Gómez, 2018)—a parameter-free, continuous dynamical system on edges—and show that it distinguishes the prism graph from $K_{3,3}$, a pair that 1-WL provably cannot separate. We then generalize it to Motif-DRESS, which replaces triangle neighborhoods with arbitrary structural motifs and converges to a unique fixed point under three sufficient conditions, and further to Generalized-DRESS, an abstract template parameterized by the choice of neighborhood operator, aggregation function and norm. Finally, we introduce $\Delta$-DRESS, which runs DRESS on each node-deleted subgraph $G\setminus\{v\}$, connecting the framework to the Kelly–Ulam reconstruction conjecture. Both Motif-DRESS and $\Delta$-DRESS empirically distinguish Strongly Regular Graphs (SRGs)—such as the Rook and Shrikhande graphs—that confound 3-WL. Our results establish the DRESS family as a highly scalable framework that empirically surpasses both 1-WL and 3-WL on well-known benchmark graphs, without the prohibitive $\mathcal{O}(n^{4})$ computational cost.
1 Introduction
The graph isomorphism problem and structural similarity measurement rely heavily on the Weisfeiler-Lehman (WL) hierarchy [7]. While the standard 1-WL test—shown to characterize the expressiveness of message-passing neural networks [8, 6]—is highly efficient, it fails to distinguish many non-isomorphic graphs. Higher-order $k$-WL tests [2] offer greater expressiveness, but their $\mathcal{O}(n^{3})$ or $\mathcal{O}(n^{4})$ complexity renders them impractical for large-scale network analysis.
The Original-DRESS algorithm [3] provides a continuous alternative: a parameter-free, non-linear dynamical system that iteratively refines edge values based on common-neighbour aggregation. As we show in Section 3, Original-DRESS distinguishes graphs that 1-WL cannot. However, it still fails to distinguish highly symmetric structures such as Strongly Regular Graphs (SRGs).
In this work, we present a hierarchy of continuous structural refinement algorithms, building from the concrete to the general:
  1. 1.
    Original-DRESS (Section 2): The foundational equation, a parameter-free dynamical system on edges.
  2. 2.
    Original-DRESS vs. 1-WL (Section 3): We show that Original-DRESS distinguishes graphs that 1-WL cannot, including the prism graph from $K_{3,3}$.
  3. 3.
    Motif-DRESS (Section 4): A generalization that replaces triangle neighborhoods with arbitrary structural motifs. Original-DRESS is recovered as a special case with $M=K_{3}$.
  4. 4.
    Generalized-DRESS (Section 5): The most abstract template, opening the aggregation function and norm as additional free parameters.
  5. 5.
    $\Delta$-DRESS (Section 6): Runs DRESS on each node-deleted subgraph $G\setminus\{v\}$, connecting the framework to the reconstruction conjecture and empirically distinguishing SRGs that confound 3-WL.
We show that Original-DRESS distinguishes graphs that 1-WL cannot, and demonstrate empirically that Motif-DRESS and $\Delta$-DRESS distinguish graphs beyond the reach of 3-WL, while maintaining scalability well below $\mathcal{O}(n^{3})$ for sparse graphs.
Isomorphism Test.
Throughout this paper, the graph fingerprint is the sorted vector of converged edge values $\operatorname{sort}(d^{*})$. Two graphs are declared non-isomorphic if and only if their fingerprints differ element-wise beyond a numerical tolerance $\epsilon$. This applies uniformly to all DRESS variants.
All experiments use a convergence tolerance of $\epsilon=10^{-6}$ and a maximum of 20 iterations; in every case convergence was reached before the iteration limit.
Related Work.
Higher-order GNN architectures that exceed 1-WL expressiveness include $k$-GNN [6], GIN [8], PPGN [5], and Ring-GNN [4]. Subgraph-based approaches such as GNN-AK+ [9] and ESAN [1] achieve higher expressiveness by running message-passing networks on node-deleted or node-marked subgraphs; $\Delta$-DRESS (Section 6) shares the node-deletion strategy but replaces learned message passing with deterministic fixed-point iteration. All of the above are supervised methods that require labeled training data and learnable parameters. In contrast, DRESS is a deterministic, unsupervised framework that produces canonical fingerprints via fixed-point iteration, without any learning or parameter tuning.
2 Original-DRESS
The Original-DRESS equation [3] is a parameter-free, non-linear dynamical system that iteratively refines edge values based on common-neighbour aggregation:
$d_{uv}^{(t+1)}=\frac{\sum_{x\in N[u]\cap N[v]}\bigl(d_{ux}^{(t)}+d_{xv}^{(t)}\bigr)}{\|u\|^{(t)}\cdot\|v\|^{(t)}}$ (1)
where $N[u]=N(u)\cup\{u\}$ denotes the closed neighborhood of $u$, and $\|u\|=\sqrt{\sum_{x\in N[u]}d_{ux}^{(t)}}$ is the node norm. The self-similarity $d_{uu}=2$ is invariant under iteration. This equation converges to a unique fixed point, providing a continuous structural fingerprint for the graph.
Self-loops. Self-loops are added to every node before iteration (i.e., the algorithm uses the closed neighborhood $N[u]=N(u)\cup\{u\}$). The self-loop edge $(u,u)$ participates in both the aggregation and the node norm; without it, an isolated edge with no common neighbours would produce $\|u\|\cdot\|v\|=0$, making the iteration undefined.
Initialization. The equation is scale-invariant (degree-0 homogeneous), so the fixed point is independent of the initial scale. Any uniform initialisation $d^{(0)}_{uv}=c$ for all edges, with $c>0$, converges to the same unique fixed point $d^{*}$. In practice, $c=1$ is used.
However, because Original-DRESS aggregates strictly over triangles (common neighbours), it cannot distinguish graphs with identical triangle counts per edge (e.g., Strongly Regular Graphs).
3 Original-DRESS Distinguishes Graphs That 1-WL Cannot
We now demonstrate a key expressiveness result: Original-DRESS—the simplest member of the DRESS family—already distinguishes graphs that 1-WL (color refinement) cannot.
Theorem 1 (DRESS distinguishes beyond 1-WL).
There exist graph pairs that 1-WL cannot distinguish but Original-DRESS can. In particular, DRESS distinguishes the prism graph ($C_{3}\square K_{2}$) from the complete bipartite graph $K_{3,3}$, a pair that 1-WL provably cannot separate.
Proof.
Both graphs are 3-regular on 6 vertices with 9 edges. Since all vertices have the same degree, 1-WL assigns a uniform color to every vertex in both graphs after every iteration, and therefore cannot distinguish them [2].
DRESS operates on edges and aggregates over common neighbours (triangles). We show that the prism must have at least two distinct edge values at convergence, while $K_{3,3}$ has a single uniform value.
$K_{3,3}$: No two adjacent vertices share a common neighbour ($K_{3,3}$ is triangle-free). Hence every edge has the same neighborhood structure, and by symmetry the unique fixed point assigns the same value to all 9 edges. Numerically, $d^{*}_{K_{3,3}}\approx 1.155$ for all edges.
Prism graph: There are two structurally distinct edge types: triangle edges (6 edges forming the two triangular faces; each pair of endpoints shares 1 common neighbour) and matching edges (3 edges connecting corresponding vertices of the two triangles; each pair of endpoints shares 0 common neighbours besides self-loops). Suppose for contradiction that all 9 edges converge to the same value $d^{*}$. Then:
  • Triangle edges: $d^{*}=\dfrac{4+4d^{*}}{2+3d^{*}}$, which gives $d^{*}=\frac{1+\sqrt{13}}{3}\approx 1.535$.
  • Matching edges: $d^{*}=\dfrac{4+2d^{*}}{2+3d^{*}}$, which gives $d^{*}=\frac{2}{\sqrt{3}}\approx 1.155$.
These are contradictory, so the prism cannot have a uniform fixed point. The converged unique edge values are:
Graph Unique values Multiplicities
Prism $\{0.922,\;1.709\}$ 3 matching edges, 6 triangle edges
$K_{3,3}$ $\{2/\sqrt{3}\approx 1.155\}$ all 9 edges identical
Since 1-WL cannot distinguish them and DRESS can, DRESS distinguishes beyond 1-WL on this instance. ∎
Remark.
This result holds for Original-DRESS alone—the simplest member of the DRESS family, using only triangle neighborhoods and $\mathcal{O}(E)$ computation per iteration. The generalizations that follow (Motif-DRESS, $\Delta$-DRESS) extend this advantage further to graphs that resist even 3-WL.
4 Motif-DRESS
Original-DRESS is limited to triangle neighborhoods. What if we use other structural motifs? Motif-DRESS generalizes Original-DRESS by defining $\mathcal{N}(u,v)$ as the set of edges that co-occur with $(u,v)$ in a specific structural motif $M$, with $f=\text{sum}$, $g=\text{geometric norm}$, and an optional weight function $w(e)$:
$d_{uv}^{(t+1)}=\frac{\displaystyle\sum_{e^{\prime}\in\mathcal{N}_{M}(u,v)}w_{e^{\prime}}\,d_{e^{\prime}}^{(t)}}{\|u\|^{(t)}\cdot\|v\|^{(t)}}$ (2)
where $\|u\|^{(t)}=\sqrt{\sum_{e^{\prime}\in\mathcal{N}_{M}(u,u)}w_{e^{\prime}}\,d_{e^{\prime}}^{(t)}}$ is the weighted node norm at time $t$.
The weights $w(e)$ act as a multiplicative factor, controlling how much structural information flows along each edge. Because the weights appear identically in the numerator and denominator (both are degree-1 in $w\cdot d$), uniformly scaling all weights does not change the fixed point; only the relative weights matter.
Different choices of motif $M$ yield different neighborhood operators:
  • Triangle ($M=K_{3}$): $\mathcal{N}(u,v)=\{(u,x),(x,v)\mid x\in N[u]\cap N[v]\}$—this recovers Original-DRESS.
  • 4-cycle ($M=C_{4}$): Aggregates over edges participating in 4-cycles with $(u,v)$.
  • $K_{4}$ clique: Aggregates over edges forming a $K_{4}$ with $(u,v)$.
4.1 Properties
Complexity: $\mathcal{O}(\text{Motif Extraction})+\mathcal{O}\!\bigl(I\cdot\sum_{e}|\mathcal{N}_{M}(e)|\bigr)$, where $I$ is the number of iterations and $\sum_{e}|\mathcal{N}_{M}(e)|$ is the total size of all motif-neighborhood lists. For Original-DRESS (triangles), this reduces to $\mathcal{O}(I\cdot m\cdot\Delta)$. For motifs such as 4-cycles or $K_{4}$ cliques in sparse graphs, the motif-neighborhood lists can be significantly smaller, making Motif-DRESS faster than Original-DRESS.
Invariant: $d_{uu}=2$. The self-similarity is invariant under iteration: since $\mathcal{N}_{M}(u,u)$ includes all edges incident to $u$ symmetrically, and $d_{ux}=d_{xu}$, substituting $v=u$ yields $d_{uu}^{(t+1)}=2\|u\|^{2}/\|u\|^{2}=2$ for all $t$. Thus $d_{uu}=2$ is a fixed property of the equation, not a free parameter.
4.2 Expressiveness
All experiments below use the $K_{4}$ clique motif. The specific SRG pairs tested below are known to be indistinguishable by 3-WL; each successful distinction therefore demonstrates that Motif-DRESS empirically exceeds 3-WL on these instances.
  • Rook vs. Shrikhande: Successfully distinguishes this pair of SRGs with parameters $(16,6,2,2)$. The Rook graph ($K_{4}\square K_{4}$) contains $K_{4}$ cliques while the Shrikhande graph does not, so the $K_{4}$-neighborhood sizes differ per edge.
  • Chang Graphs: Distinguishes all 6 pairwise comparisons among the four SRGs with parameters $(28,12,6,4)$: T(8), Chang-1, Chang-2, and Chang-3.
4.3 Proof Sketch of Convergence
Theorem 2 (Convergence of Motif-DRESS).
Let $F:\mathbb{R}_{+}^{|E|}\to\mathbb{R}_{+}^{|E|}$ be the Motif-DRESS update operator. If the aggregation function and the node norm satisfy the degree-0 homogeneity condition ($p=2q$), then $F$ converges to a unique fixed point $d^{*}\in[0,2]^{|E|}$ for any non-negative initial state $d^{(0)}>0$.
Proof sketch.
The argument proceeds in three steps:
  1. 1.
    Scale Invariance (Degree-0 Homogeneity): The aggregation function is positively homogeneous of degree $p$, and the norm of degree $q$. Since $p=2q$, the ratio $F(d)$ is homogeneous of degree 0: for any $\lambda>0$, $F(\lambda d)=F(d)$. Hence the iteration is self-regularizing and cannot diverge or collapse to zero.
  2. 2.
    Boundedness: For any valid symmetric neighborhood operator $\mathcal{N}_{M}(e)$, the Cauchy–Schwarz inequality (or Hölder’s inequality for Minkowski-$r$ variants) bounds the numerator by the denominator. Self-loop inclusion ensures the denominator is strictly positive. Thus $F(d)_{uv}\in[0,2]$ for all $d>0$.
  3. 3.
    Contraction on the Hilbert Projective Metric: Since $F$ is a positive, degree-0 homogeneous map on the cone $\mathbb{R}_{>0}^{|E|}$, Birkhoff’s contraction theorem guarantees that $F$ is a strict contraction under the Hilbert projective metric $d_{H}(x,y)=\log\!\bigl(\max_{e}x_{e}/y_{e}\cdot\max_{e}y_{e}/x_{e}\bigr)$, provided $F$ maps a bounded part of the cone into a strictly smaller part—which follows from the Cauchy–Schwarz bound above. By the Banach fixed-point theorem on the complete metric space $(\mathbb{R}_{>0}^{|E|}/\sim,\,d_{H})$, the iteration converges to a unique ray, and the boundedness step pins the representative to $d^{*}\in[0,2]^{|E|}$.
A complete formal verification of the contraction constant is deferred to future work; all empirical tests confirm convergence within 20 iterations. ∎
5 Generalized-DRESS
Motif-DRESS fixes the aggregation to summation and the norm to the geometric mean. Generalized-DRESS is the most abstract template, allowing any choice of these components. For each edge $(u,v)$ and a symmetric neighborhood operator $\mathcal{N}(u,v)$:
$d_{uv}^{(t+1)}=\frac{f\bigl(\{d_{e^{\prime}}^{(t)}\}_{e^{\prime}\in\mathcal{N}(u,v)}\bigr)}{g^{(t)}(u)\cdot g^{(t)}(v)}$ (3)
where:
  • $\mathcal{N}(u,v)$ is a symmetric neighborhood operator—the set of edges aggregated for edge $(u,v)$,
  • $f$ is the aggregation function over edges in $\mathcal{N}(u,v)$,
  • $g(u)$ is the node norm.
Original-DRESS and Motif-DRESS are both special cases: Original-DRESS fixes $\mathcal{N}$ to triangles, $f=\text{sum}$, $g=\text{geometric norm}$; Motif-DRESS generalizes $\mathcal{N}$ to arbitrary motifs while keeping the same $f$ and $g$. Generalized-DRESS opens all three parameters, enabling variants such as Cosine-DRESS (cosine similarity aggregation) or Minkowski-$r$ norms.
6 $\Delta$-DRESS
$\Delta$-DRESS breaks symmetry by running DRESS on each node-deleted subgraph $G\setminus\{v\}$ for every $v\in V$. The graph fingerprint is the sorted multiset of $n$ converged edge-value vectors (one per deletion), compared by flattening and sorting without any summarization.
Deleting a node from a regular graph produces an irregular subgraph where DRESS can now distinguish structure that was hidden by the uniform regularity.
6.1 Connection to the Reconstruction Conjecture
The multiset $\{\!\{\text{DRESS}(G\setminus\{v\}):v\in V\}\!\}$ is directly analogous to the deck in the Kelly–Ulam reconstruction conjecture, which posits that graphs with $n\geq 3$ are determined (up to isomorphism) by their multiset of node-deleted subgraphs. $\Delta$-DRESS computes a continuous relaxation of this deck.
6.2 Properties
Complexity: $\mathcal{O}(n\cdot I\cdot m\cdot\Delta)$, where $I$ is the number of iterations per deletion. The computation for each deleted subgraph is entirely independent, making $\Delta$-DRESS embarrassingly parallel.
6.3 Expressiveness
$\Delta$-DRESS empirically distinguishes the following non-isomorphic pairs:
  • Rook vs. Shrikhande: Successfully distinguished (SRG(16,6,2,2)(16,6,2,2); confounds 3-WL).
  • Chang Graphs: Distinguished all 6 pairs among T(8) and the three Chang graphs (SRG(28,12,6,4)(28,12,6,4); confound 3-WL).
  • $2\times C_{4}$ vs. $C_{8}$: Successfully distinguished (both 2-regular on 8 nodes).
  • Petersen vs. Pentagonal Prism: Successfully distinguished (both 3-regular on 10 nodes).
7 Family Structure
The DRESS variants introduced in this paper form a nested hierarchy with an orthogonal composition operator:
$\text{Generalized-DRESS}\supset\text{Motif-DRESS}\supset\text{Original-DRESS}$
$\Delta(\cdot)$ is an orthogonal wrapper applicable to any of the above: given any DRESS variant $\mathcal{F}$,
$\Delta\text{-DRESS}(\mathcal{F})=\{\!\{\mathcal{F}(G\setminus\{v\}):v\in V\}\!\}$
The node-deletion strategy is independent of the choice of $\mathcal{N}$, $f$, and $g$. All experiments in Section 6 use $\mathcal{F}=\text{Original-DRESS}$, but $\Delta$-Motif-DRESS or $\Delta$-Cosine-DRESS are equally valid and may offer complementary expressiveness. The deletion strategy generalizes naturally to $\Delta^{k}$-DRESS, which runs DRESS on all $\binom{n}{k}$ node-deleted subgraphs; we leave the exploration of this hierarchy to future work.
8 Conclusion
We presented the DRESS family, a hierarchy of continuous structural refinement algorithms. Starting from Original-DRESS—a single equation proposed in 2018—we showed it distinguishes graphs that 1-WL cannot (Theorem 1). We then generalized it to Motif-DRESS, which replaces triangle neighborhoods with arbitrary structural motifs and converges to a unique fixed point under three sufficient conditions. Generalized-DRESS abstracts further, opening the aggregation function and norm as free parameters. $\Delta$-DRESS extends the framework by running DRESS on each node-deleted subgraph, connecting it to the Kelly–Ulam reconstruction conjecture and empirically distinguishing Strongly Regular Graphs that confound 3-WL, while maintaining scalability well below $\mathcal{O}(n^{3})$ for sparse graphs. These results establish the DRESS family as a highly scalable framework that empirically surpasses both 1-WL and 3-WL, without the prohibitive computational cost of higher-order WL tests. An open-source implementation with bindings for C, C++, Python, Rust, Go, Julia, R, MATLAB, and WebAssembly is publicly available at https://github.com/velicast/dress-graph.